after the collision. 10. A hoop and a disk are allowed to roll (without slipping) down an incline plane. If they both start from rest, from one meter above the bottom of the plane, how fast is each object traveling at the bottom of the plane? h 1m
Answer #1

acceleation of rolling body can be defined by

a = g\sin\theta/(1+K^2/R^2)

where MK^2 = I = moment of inertia of the body

sinceg, theta and M are fixed and Moment of inertia of hoop is MR^2 while that of solid cylinder is MR^2/2. So putting K^2/R^2 = 1/2 for solid cylinder and 1 for hoop

we will get acclnof solid cylinder is greater than that of hoop